# Timestamp Manipulation: Timestamp-based Nakamoto-style Blockchains are Vulnerable

### v4

@@ IX Conclusion
Our simulations validate SUUM’s dominance over prior attacks. Crucially, SUUM bypasses traditional constraints like hash power thresholds by exploiting timestamp calibration, maintaining minimal difficulty escalation despite sustained exploitation. This asymmetric advantage stems from the protocol’s inability to self-correct reward distortions, leading to a death spiral where adversarial coalitions dominate and honest participation becomes economically untenable.
@@ Appendix A Related Work
We now embark on a discussion concerning the related work of this paper, primarily focusing on three aspects: attacks targeting vulnerabilities in incentive models and consensus protocols of Nakamoto-like blockchains, Ethereum 2.x-like blockchains, and other facets.
@@ A-A Attacks on Incentive Model and Consensus Protocol of Nakamoto-like Blockchains
Withholding Attack. In Nakamoto-like blockchains, Withholding Attack represents a prevalent malicious behavior [30, 37, 38, 39, 40, 41, 42, 43]. Most consensus-layer attacks rely on block withholding and specific capabilities of the adversary, exemplified by the Selfish Mining attack, where the adversary withholds blocks and strategically chooses the timing of their release to increase their share of blocks. Variants such as Stubborn Mining attacks involve the adversary continuing to mine even when their selfish fork lags behind the honest chain [38, 44]. These attacks undermine the fairness and normal operation of blockchain networks. By impeding the timely dissemination of certain blocks, adversaries attempt to control the pace of block generation within the network and disrupt the normal functioning of honest nodes. This not only results in the potential waste of computing power and resources of honest nodes but also complicates and destabilizes the network’s fork situation, reducing its security and reliability. Consequently, user trust in Nakamoto-like blockchains is eroded, exerting a negative impact on the entire blockchain ecosystem.
Bribing Attack. In Nakamoto-like blockchains, the Bribing Attack poses a severe threat to network security and stability [45, 46, 47, 48, 49]. First described by Bonneau, this type of attack encompasses multiple mechanisms. Early research indicated that bribery attacks may intensify when participants rely solely on transaction fees for compensation. For instance, bribery attacks implemented through smart contracts exhibit inbound and outbound modes, with subsequent research continuously enhancing their efficiency, such as through cross-system double-spending conspiracies or leveraging smart contract payments. The motives for these attacks are diverse, including double-spending, consensus disruption, censorship, transaction front-running, or the sabotage of hash-time locked contracts (HTLCs) [50, 51]. Recent research has proposed bribery attack models targeting longest-chain blockchains (such as Bitcoin), with some assuming a dichotomy between honest and rationally bribe-accepting participants, employing inbound bribery (paid in native tokens). Other models may be more general, considering rationality on all sides, with outbound bribery (e.g., paid in USD) unaffected by token price fluctuations. These attacks significantly disrupt the normal operational order of blockchains.
DoS attack. In Nakamoto-like blockchains, Denial of Service (DoS) attacks primarily disrupt system availability by consuming network resources [52, 53, 54, 55]. At the network layer, adversaries often launch traffic flooding attacks, such as sending a massive number of false requests or packets to target nodes, rapidly consuming network bandwidth and causing nodes to become overwhelmed with processing these invalid flows, thereby preventing them from timely responding to legitimate transactions. For example, a deluge of spam transaction requests can clog network channels, hindering the smooth propagation and processing of normal transactions. Such attacks significantly interfere with the normal operation of blockchain networks, increasing transaction confirmation times and reducing system efficiency. They undermine the decentralized feature of blockchain, as nodes under attack are unable to participate equally in network activities. Furthermore, DoS attacks may trigger a chain reaction, affecting users’ trust in blockchain and impeding its application and development across various domains.
Timestamp Attack. In Nakamoto-like blockchains, Timestamp Manipulation Attacks exploit the configurable feature of block timestamps, allowing adversaries to meticulously set timestamps that may interfere with the normal generation and confirmation processes of blocks [16]. For instance, adversaries can manipulate mining difficulty adjustments by setting unreasonable timestamps, thereby impacting their mining rewards or network stability. This attack vector potentially undermines the fairness and normal operational order of blockchain, granting adversaries undue advantages and posing a latent threat to the security and reliability of Nakamoto-like blockchains. Some research has addressed the role of timestamps in attacks, such as in Ethereum where timestamps were used to attack smart contracts. However, there is limited research on timestamp attacks at the consensus layer. Stretch identified two timestamp vulnerabilities in the geth codebase, one of which could be used to reduce honest mining difficulty, but did not delve into the impact on consensus mechanisms.
It is worth mention that the UUM attack proposed in our paper falls within the category of timestamp attacks, while SUUM combines withholding attacks with timestamp attacks.
@@
Withholding Attack. In Ethereum 2.x-like PoS blockchains, the Withholding Attack poses a threat to network stability. An adversary may withhold blocks to prevent honest nodes from propagating them, releasing the withheld blocks at an opportune moment to gain undue benefits [56, 57, 58, 59]. This attack resembles the withholding attack observed in Nakamoto-style blockchains but exhibits distinct characteristics due to the unique protocol mechanisms of Ethereum 2.x PoS. For instance, adversaries can exploit the staking and voting mechanisms of validator nodes to disrupt the normal block confirmation process during consensus. By withholding blocks they have mined or validated, they can influence the block height and fork conditions within the network, thereby undermining network consistency and certainty. Such attacks may result in wasted efforts by honest nodes, increase network uncertainty, and diminish user trust in the network. Furthermore, they may afford the adversary an advantageous position in subsequent block competitions or network decisions, thereby impacting the fairness and security of the entire Ethereum 2.x-like blockchain network.
Nothing-at-stake Attack. In Ethereum 2.x-like PoS blockchains, the Nothing-at-Stake Attack represents a severe threat to network activity and performance [60, 61, 62]. Due to the characteristics of the staking distribution and validation mechanisms within the PoS system, adversaries can exploit this vulnerability to contribute effortlessly to multiple forks. Without bearing any actual risk, adversaries can interfere with the normal transaction confirmation and block production processes by operating on different forks. This leads to a chaotic situation of network forks, making it difficult for the system to achieve a unified consensus and reducing the network’s efficiency in processing transactions. For example, under normal circumstances, validators should choose to support a single correct fork based on their own staking and network rules. However, adversaries can freely stake on multiple forks, disrupting the stability and reliability of the network. This, in turn, affects users’ trust in the Ethereum 2.x-like blockchain network, hinders its normal development, and has a negative impact on the entire blockchain ecosystem.
Long-range Attack. The Long-Range Attack is a potential threat to Ethereum 2.x-like PoS blockchains, exploiting a vulnerability in the PoS mechanism where adversaries can secretly construct an alternative chain starting from an earlier point in the blockchain history [63, 61, 64]. This attack is named for the adversaries’ ability to backtrack to a long-range past point in the blockchain without significant resource consumption and initiate a new chain from there. If successful, the previous transaction records and states of the network will become untrusted, undermining the consistency principle of blockchain data. For instance, adversaries can alter transaction information in past blocks, reverse completed transaction states, or create new transactions out of thin air, severely disrupting the normal operational order of the network. This attack leverages potential flaws in Ethereum 2.x-like PoS mechanism in managing historical data, dealing a severe blow to the trust foundation of the entire blockchain ecosystem. It may spark user concerns about network security, thereby impacting the widespread adoption and sustainable development of Ethereum 2.x-like PoS blockchains. Consequently, continuous enhancements in preventive measures are necessary to defend against such attacks.
DoS Attack. In Ethereum 2.x-like PoS blockchains, DoS attacks manifest in various forms and pose severe threats. At the network layer, DoS attacks resemble traditional patterns, where adversaries launch flood attacks by sending a vast number of fake requests or data packets to occupy network bandwidth and node resources, thereby preventing legitimate transactions from being processed normally. Attacks at the application layer primarily target smart contracts and transaction pools (Txpool) [65, 66, 67]. At the smart contract level, adversaries exploit vulnerabilities or design flaws in contracts to initiate malicious transactions that cause contract execution to enter infinite loops or exhaust computational resources. For example, triggering complex logic operations without resource limits can lead to contract execution timeouts and node resource depletion, resulting in a DoS condition that affects the normal operation of the network. In terms of the Txpool, adversaries send a large number of transactions with low gas prices but high computational complexity, causing congestion in the Txpool and preventing normal transactions from being processed in a timely manner. This can also manipulate the order of transactions, influence participants’ choices in packaging transactions, interfere with the normal entry of transactions into the blockchain network, and undermine the fairness and efficiency of transaction processing in the network.
Reorg Attack. The Reorg Attack in Ethereum 2.x-like PoS blockchains poses a severe threat to network security and stability [56, 68]. This attack increases the proportion of blocks belonging to Byzantine validators on the main chain to gain more profits or downgrade chain quality or performance. It encompasses Short Reorg Attacks (SRA) and Long Reorg Attacks (LRA). By leveraging controlled validator nodes, adversaries release new block branches at specific times, causing the original main chain to be replaced. This results in the transaction states on the previous main chain becoming uncertain, affecting user fund security and transaction reliability, disrupting block order and transaction confirmation processes, and weakening users’ trust in the network.
A-C Attacks on Other Aspects
Smart Contract Attack. In Ethereum 2.x-like PoS blockchains, the attack vectors targeting smart contracts are diverse and pose significant risks [69, 70, 71]. Reentrancy vulnerabilities represent one of the common attack methods, where adversaries exploit flaws in the contract’s failure to correctly update its state during external calls, resulting in repeated invocations of contract functions and the abnormal transfer of contract assets. The infamous DAO incident suffered substantial losses due to a reentrancy vulnerability. Overflow vulnerabilities are also noteworthy, as errors in computation may arise when data exceeds the range of variable types, disrupting the normal logical execution of contracts. Furthermore, adversaries exploit vulnerabilities in contract permission management to obtain unauthorized operational privileges, allowing them to tamper with contract data or execute malicious operations. These attacks not only render the contracts ineffective in fulfilling their intended functions but may also lead to user fund losses, transaction data chaos, and severely impact the stability and trust of the Ethereum ecosystem. While formal verification and fuzz testing techniques offer certain defenses to a degree, new attack vectors continue to emerge, necessitating ongoing research and the enhancement of defensive measures.
Txpool Attack. In Ethereum 2.x-like PoS blockchains, attacks against the Txpool primarily exploit its mechanism characteristics to disrupt transaction processing [55, 66]. Adversaries often employ a strategy of submitting a large number of transactions with low gas prices but high computational complexity, thereby causing congestion within the Txpool. Since the Txpool typically sorts and processes transactions based on factors such as gas price, these malicious transactions, despite paying less gas, consume substantial resources for verification and processing due to their computational complexity. This leads to a situation where legitimate transactions cannot be timely packaged by participants, thereby extending transaction confirmation times. Furthermore, adversaries may attempt to manipulate the order of transactions within the Txpool by controlling the timing of transaction broadcasts and their queuing positions, thereby influencing participants’ selection of transactions. This enables certain specific transactions to be prioritized or delayed in processing, disrupting the fairness and normal flow of transaction processing. Consequently, this affects the transaction efficiency and user experience of the entire Ethereum 2,0-like blockchain network, posing a threat to the network’s normal operation.
Data Storage Attack. In Ethereum 2.x-like PoS blockchains, attacks targeting data storage pose significant threats [72]. Adversaries may exploit vulnerabilities in the storage mechanism to illegally manipulate storage pointers, maliciously tampering with account balances, data, or access control information, resulting in erroneous or inconsistent account states. Additionally, adversaries can leverage the characteristics of the data structures used in account storage by carefully crafting accounts that increase the cost of account access upon insertion, thereby achieving the purpose of raising storage and access cost. For instance, modifying account balance data can lead to abnormal fund displays or illegal transfers; disrupting access controls can grant unauthorized account operation privileges. Such attacks may also interfere with the account state update process, causing account states to become disconnected from actual transaction conditions. If successful, these attacks can lead to severe consequences such as user asset losses and transaction record chaos, undermining the accurate processing and trust mechanisms of account information within the entire blockchain system.

### v5

@@ IX Conclusion
Our simulations validate SUUM’s dominance over prior attacks. Crucially, SUUM bypasses traditional constraints like hash power thresholds by exploiting timestamp calibration, maintaining minimal difficulty escalation despite sustained exploitation. This asymmetric advantage stems from the protocol’s inability to self-correct reward distortions, leading to a death spiral where adversarial coalitions dominate and honest participation becomes economically untenable. We have exposed four mining pools that conduct timestamp manipulation attacks in three major Ethereum 1.x-based blockchain systems.
@@ Appendix A Related Work
We now embark on a discussion concerning the related work of this paper, primarily focusing on three aspects: attacks targeting vulnerabilities in incentive models and consensus protocols of Nakamoto-like and Ethereum 2.x-like blockchains.
@@ A-A Attacks on Incentive Model and Consensus Protocol of Nakamoto-like Blockchains
Withholding Attack. Withholding attacks are prevalent malicious behaviors in Nakamoto-like blockchains [25, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Most consensus-layer attacks (e.g., selfish mining, stubborn mining) rely on block withholding: adversaries withhold blocks to control release timing or continue mining on lagging selfish forks [34, 27], undermining network fairness, wasting honest nodes’ computing resources, and reducing network security and reliability.
Bribing Attack. Bribing attacks pose severe threats to the security and stability of Nakamoto-like blockchains [42, 43, 44, 45, 46, 47, 48]. First proposed by Bonneau, these attacks include multiple mechanisms (e.g., inbound/outbound bribery via smart contracts). Their motives range from double-spending and consensus disruption to censorship and HTLC sabotage [49, 50], significantly disrupting blockchain operations.
DoS attack. Denial of Service (DoS) attacks disrupt system availability by consuming network resources [51, 52, 53, 54]. At the network layer, adversaries often launch traffic flooding attacks (e.g., fake requests) to exhaust bandwidth, hindering legitimate transaction processing, increasing confirmation latency, and weakening blockchain decentralization.
Timestamp Attack. Timestamp manipulation Attacks exploit configurable block timestamps in Nakamoto-like blockchains [14], interfering with block generation/confirmation or manipulating difficulty adjustments. Research on timestamp attacks at the consensus layer remains limited: while Stretch identified timestamp vulnerabilities in Geth (e.g., reducing honest mining difficulty), the impact on consensus mechanisms was not explored. Note that the UUM attack proposed in this paper belongs to timestamp attacks, and SUUM combines withholding attacks with timestamp attacks.
@@
Withholding Attack. In Ethereum 2.x-like PoS blockchains, withholding attacks threaten network stability [55, 56, 57, 58]. Adversaries withhold blocks to block propagation and release them opportunistically, exploiting staking/voting mechanisms to disrupt block confirmation, affect block height/forks, and undermine network consistency.
Nothing-at-stake Attack. Nothing-at-stake attacks severely threaten the activity and performance of Ethereum 2.x-like PoS blockchains [59, 60, 61]. Leveraging PoS staking/validation flaws, adversaries participate in multiple forks without risk, causing fork chaos, hindering consensus unification, and reducing transaction processing efficiency.
Long-range Attack. Long-range attacks are potential threats to Ethereum 2.x-like PoS blockchains [62, 60, 63]. Adversaries secretly construct alternative chains from historical blockchain points (with low resource consumption), undermining data consistency (e.g., altering past transactions) and eroding the trust foundation of PoS blockchains.
DoS Attack. DoS attacks on Ethereum 2.x-like PoS blockchains take multiple forms [64, 65, 66]. Network-layer attacks involve traffic flooding, while application-layer attacks target smart contracts (e.g., infinite loops) and Txpool (e.g., low-gas high-complexity transactions), causing resource exhaustion or congestion.
Reorg Attack. Reorg attacks threaten the security and stability of Ethereum 2.x-like PoS blockchains [55, 67]. Adversaries increase the proportion of Byzantine validator blocks via controlled nodes (including Short/Long Reorg Attacks), replacing the main chain, making transaction states uncertain, and weakening user trust.